# Diffusion Models as Dataset Distillation Priors

### v1

@@ A.4.4 Cross-dataset generalization
We posit that the cross-datasets evaluation is essential to measure the generalization and versatility of a DD method and its datasets, which is overlooked by most DD methods. According to Su et al. (2024), we extract 200 categories, which are predefined in Le & Yang (2015), from the ImageNet-1K dataset distilled by DAP as the distilled Tiny-ImageNet dataset. Table 9 shows that the extracted subsets (end with “-G”) maintain strong validation performance on the target set, thereby confirming that our distilled data not only preserves the utility of the original dataset but also supports effective reuse across datasets. The results highlight the advantage of DAP: the distilled dataset is not tied to a single dataset domain but can be flexibly transferred and reused.
A.4.5 Early Stop strategy
In the guided sampling process, we employ the early stop guidance mechanism, which enhances sampling quality by only guiding earlier diffusion timesteps than the entire timesteps, thereby providing a better trade-off between sample diversity and fidelity (Chen et al., 2025; Chan-Santiago et al., 2025). Besides, applying representativeness prior guidance in the early stage of the sampling trajectory also reduces the sampling cost (refer to section A.4.6). We summarize the sampling process with an early stop strategy in algorithm 2. To evaluate its effectiveness, we conducted experiments with different stopping parameters $t_{stop}$. The mechanism deactivates guidance for timesteps $t<t_{stop}$ in the reverse process, $t_{stop}=0$ means complete guidance, while $t_{stop}=50$ represents no guidance. The qualitative and quantitative results, as illustrated in fig. 8, indicate that $t_{stop}=25$ yields the best performance.
@@ A.4.5 Early Stop strategy
A.4.6 Sampling-time scaling
@@
A.5 Visualizations
A.5.1 Representativeness guidance scale

### v2

@@ A.3.5 Cross-dataset generalization
We posit that the cross-datasets evaluation is essential to measure the generalization and versatility of a DD method, which is overlooked by most DD methods. According to Su et al. (2024), we extract 200 categories, which are predefined in Le and Yang (2015), from the ImageNet-1K dataset distilled by DAP as the distilled Tiny-ImageNet dataset. Table 10 shows that the extracted subsets (end with “-G”) maintain strong validation performance on the target set, thereby confirming that our distilled data not only preserves the utility of the original dataset but also supports effective reuse across datasets. The results highlight the advantage of DAP: the distilled dataset is not tied to a single dataset domain but can be flexibly transferred and reused.
A.3.6 Early Stop strategy
In the guided sampling process, we employ the early stop guidance mechanism, which enhances sampling quality by only guiding earlier diffusion timesteps than the entire timesteps, thereby providing a better trade-off between sample diversity and fidelity (Chen et al., 2025; Santiago et al., 2025). Besides, applying representativeness prior guidance in the early stage of the sampling trajectory also reduces the sampling cost (refer to section A.3.7). We summarize the sampling process with an early stop strategy in algorithm 2. To evaluate its effectiveness, we conducted experiments with different stopping parameters $t_{stop}$. The mechanism deactivates guidance for timesteps $t<t_{stop}$ in the reverse process, $t_{stop}=0$ means complete guidance, while $t_{stop}=50$ represents no guidance. The qualitative and quantitative results, as illustrated in fig. 8, indicate that $t_{stop}=25$ yields the best performance.
@@ A.3.6 Early Stop strategy
A.3.7 Sampling-time scaling
@@
A.4 Visualizations
A.4.1 Representativeness guidance scale
@@ A.4.2 Representativeness comparison
A.5 Disclosure of the use of Large Language Models
Given that the use of large language models (LLMs) is allowed as a general-purpose assist tool, this work utilizes LLMs to polish the sentences of the article. There is no significant role in research ideation and writing to the extent that they cannot be regarded as contributors.